# Enhancing Arabic Automated Essay Scoring with Synthetic Data and Error Injection

### v1

@@ 5.3 Controlled Error Injection
Zaebuc-1 22.44 57.58 24.50 23.33 26.76 61.06 84.85 43.70 42.50 45.31
Zaebuc-1 + GPT essays 14.92 60.61 26.43 25.55 27.45 96.00 96.97 92.32 98.04 88.89
Zaebuc-1 + BERT errors 27.87 57.58 38.02 35.86 44.93 82.70 87.88 71.66 70.83 74.38
Zaebuc-1 + GPT errors_1 17.14 57.58 25.64 25.18 26.27 96.47 96.97 94.16 97.92 91.67
Zaebuc-1 + GPT errors_0 20.84 57.58 32.76 31.53 46.08 93.79 93.94 95.12 96.49 94.44
Table 6: Performance comparison of different training datasets. GPT essays are the original correct essays generated from GPT-4o, BERT errors are the erroneous essays using the controlled injection BERT model, GPT errors_1 are the erroneous essays using GPT-4o with one shot error example, while GPT errors_0 with Zero shot settings
6 Evaluation
6.1 Metrics
Our primary evaluation metric is Quadratic Weighted Kappa (QWK) Cohen (1968), the most widely-used metric in recent AES research Ke and Ng (2019). We also report accuracy, macro precision (P), recall (R), and F1 scores. Model predictions are evaluated against the average of the gold holistic CEFR labels and in a multi-reference setting against the three individual gold labels per essay (§3).
6.2 Data
We use the ZAEBUC dataset for all the experiments. The dataset was split (70%,15%,15%) for Train, Dev and Test sets, respectively Alhafni et al. (2023). Table 6 shows the performance results of the BERT-based Scoring Model utilizing different training sets.
@@ 6.3 Model
We treat the task of automated essay scoring (AES) as a text classification problem. For Arabic AES, we fine-tune CAMeLBERT MSA Inoue et al. (2021) on the training split of ZAEBUC. The models were trained by using the average CEFR gold labels. During training, we ignore the essays that are labeled as Unassessable, but we penalize the models for missing them in the evaluation. We fine-tune the models for 5 epochs, with a maximum sequence length of 512, a learning rate of 5e-5, and a batch size of 32.
@@ 6.4 Results
Our results are presented in Table 6. Our baseline system, only trained on the ZAEBUC training set, indicates room for improvement, with the F1 at 24.50% and QWK at 22.44%. We then switched between different datasets to enhance the Arabic AES model and to measure the impact of data augmentation on the model. We introduced the 3,040 correct essays generated using GPT-4o, which dropped the QWK compared to the baseline while increasing the F1. However, we observed significant improvements in the multi-reference settings for the QWK (96.00%) and F1 (92.32%). Since this is an error-free, synthetic dataset, we further refine the model by introducing essays that mimic human-like writing styles. We compare the two approaches introduced in section 5: (1) GPT-based error injection (with and without error instruction examples) and (2) the controlled error injection method based on CAMeLBERT MSA. Inoue et al. (2021). The results demonstrate that the controlled error model enhances performance across all metrics, particularly in the average-reference setting, that achieved 27.87 % and 38.02 for QWK and F1 respectively. However, the GPT-based error injection performs best under the multi-reference setting, where the one-shot essays increased the QWK to 96.47 and the zero-shot erroneous essays improved the F1 to 95.12%. These results suggest that LLMs and Transformer methods generalize effectively in AES models and also improve performance. The results emphasize that a controlled methodology is more reliable in mimicking patterns of learner errors and mistakes. This positions the controlled BERT-based model as a vital component of the data augmentation pipeline, effectively addressing the limitations of the GPT models.
@@ 7 Conclusions and Future Work
This paper proposes a hybrid framework for building an Arabic AES system. We utilize both LLMs and Transformers to overcome the challenges of dataset scarcity by generating synthetic data that mimics the writing styles of Arabic learners. We leverage the ZAEBUC corpus and build linguistic feature profiles as well as error profiles that align with each CEFR score. We prompt GPT-4o to generate 3,040 Arabic essays based on multiple CEFR levels corresponding to 152 prompts. However, GPT-4o’s effectiveness depends heavily on prompt engineering, achieving only 27.5% agreement with the original linguistic feature profiles. To introduce errors into GPT-generated essays, we compare two approaches: (1) GPT-4o error injection, which struggles with realistic learner writing due to misinterpreted linguistic errors or complex prompts, causing inconsistencies; (2) a controlled method fine-tuning the CAMeLBERT MSA model to inject errors proportionally to their natural occurrence. The latter method is an essential component of data augmentation for Arabic AES. Our evaluation, using a fine-tuned BERT classifier, shows that the proposed hybrid framework that combines GPT-generated data and a controlled error injection method shows consistent improvement over the baseline (QWK: 27.87%, F1: 38.02%), achieving more predictable and interpretable results. These results suggest the controlled approach to generating learner errors better captures the actual distribution of errors across proficiency levels. In future work, we plan to expand the dataset by incorporating real essay samples, which will enhance the linguistic and error profiles for each CEFR level, thereby improving the proposed data augmentation model. In addition, we will deploy the AES system as part of an interactive educational writing assistant tool to help students enhance their Arabic writing with immediate feedback on errors and CEFR scores.
8 Limitations
@@
Despite the effectiveness of our hybrid Arabic AES framework, we note several limitations related to the quality of generated Arabic essays, error injection accuracy, and the generalization of the AES model. The lack of A1 and C2 essays in ZAEBUC means that there is no gold reference data for these levels, which may impact both linguistic and error profiles, affecting the accuracy of GPT-generated essays. Furthermore, different biases exist in both ZAEBUC and the GPT outputs: ZAEBUC overrepresents B-level essays, skewing model training; and GPT-4o tends to favor Arabic culture themes, thereby limiting diversity in the generated texts, and usability in other contexts. In addition, due to the lack of comprehensive gold data, GPT struggles to fully replicate real learner writing styles, achieving only 27.5% agreement with linguistic feature profiles. Another limitation is the model’s ability to generalize across various domains and question types. The AES system may struggle with broader writing tasks and alternative prompts since the dataset and augmentation methods focus on predefined prompts. Relying solely on CEFR as a holistic scoring method limits interpretability. Enhancing the dataset with multi-trait annotations, such as coherence, argumentation, and organization, could improve scoring accuracy and feedback quality. Moreover, better-controlled GPT prompting could refine the quality and diversity of generated essays, reducing biases and improving alignment with real learner writing patterns.
9 Ethical Considerations
@@
Appendix A GPT-4o Error injection algorithm
Figure 6: GPT-4o Error injection and verification algorithm

### v2

@@ 5.3 Controlled Error Injection
ZAEBUC (baseline) 22.44 57.58 24.50 23.33 26.76 61.06 84.85 43.70 42.50 45.31
ZAEBUC + GPT essays 14.92 60.61 26.43 25.55 27.45 96.00 96.97 92.32 98.04 88.89
ZAEBUC + BERT errors 27.87 57.58 38.02 35.86 44.93 82.70 87.88 71.66 70.83 74.38
ZAEBUC + GPT errors_1 17.14 57.58 25.64 25.18 26.27 96.47 96.97 94.16 97.92 91.67
ZAEBUC + GPT errors_0 20.84 57.58 32.76 31.53 46.08 93.79 93.94 95.12 96.49 94.44
Table 6: Performance comparison of different training datasets. GPT essays are the original correct essays generated from GPT-4o, BERT errors are the erroneous essays using the controlled injection BERT model, GPT errors_1 are the erroneous essays using GPT-4o with one-shot error example, while GPT errors_0 with Zero-shot settings.
6 Experimental Setup
This study focuses on introducing a data augmentation framework and synthetic Arabic essay corpus, rather than proposing a new AES model. We use a BERT-based model trained on the original ZAEBUC dataset as the reference baseline, evaluating how different augmentation strategies (e.g., GPT-4o generation, BERT-based error injection) improve performance relative to this setup.
6.1 Data and Metrics
We use the ZAEBUC dataset for all the experiments, following the splits created by Alhafni et al. (2023): 70% Train, 15% Dev, and 15% Test. Our primary evaluation metric is Quadratic Weighted Kappa (QWK) Cohen (1968), the most widely used metric in AES research Ke and Ng (2019). We also report accuracy, macro precision (P), recall (R), and F1 scores. Model predictions are evaluated in two settings settings: average-refence and multi-reference. The average-reference setting uses the rounded average of the three scores as the gold label, while the multi-reference considers each of the three human-assigned labels as a valid reference during evaluation, following a more tolerant evaluation strategy (§3).
@@ 6.2 Model
We treat the task of AES as a text classification problem. We fine-tune CAMeLBERT MSA Inoue et al. (2021) on the training split of ZAEBUC. The models were trained by using the average CEFR gold labels. During training, we ignore the essays that are labeled as Unassessable, but we penalize the models for missing them in the evaluation. We fine-tune the models for 5 epochs, with a maximum sequence length of 512, a learning rate of 5e-5, and a batch size of 32.
@@ 6.3 Results
Our results are presented in Table 6. Our baseline system, only trained on the ZAEBUC training set, indicates room for improvement, with the F1 at 24.50% and QWK at 22.44%. We then switched between different datasets to measure the impact of data augmentation on the model.
Impact of Synthetic Data
We tested data augmentation by adding 3,040 corrected GPT-4o-generated essays, which lowered QWK but increased F1. Notably, the multi-reference setting saw significant gains, with QWK at 96.00% and F1 at 92.32%. This pattern stems from the flexibility of multi-reference evaluation, which treats all three human-assigned CEFR labels as valid references. This accommodates natural scoring variations and increases the chance that model predictions, especially on synthetic data, align with at least one reference label, boosting QWK and F1 scores for both GPT-generated and error-injected essays.
Comparison of Error Injection Methods
As the initial synthetic essays were error-free, we further refined the model by adding essays with human-like errors. We compared two methods from §5: (1) GPT-based error injection (with and without instruction examples) and (2) the controlled BERT-based method. The results demonstrate that the controlled error model improves performance in all metrics, particularly in the average reference setting, which achieved 27.87 % and 38.02 for QWK and F1, respectively. This result aligns with expectations, as the BERT-injected errors closely follow CEFR-based error distributions, producing errors that realistically reflect learner writing and better match the average of human ratings. GPT-based error injection performed best in the multi-reference setting, with one-shot examples reaching 96.47% QWK and zero-shot boosting F1 to 95.12%. While less aligned with CEFR profiles, GPT errors benefit from fluency and variability, increasing the chance of matching at least one human reference in this flexible evaluation.
7 Discussion
This study demonstrated the effectiveness of synthetic data and controlled error injection in enhancing Arabic AES, providing key insights into metric interpretation, data expansion, and methodological choices. First, we emphasize that QWK offers a more robust metric than accuracy for evaluating AES systems, particularly under imbalanced class distributions. Unlike accuracy, which is biased by majority classes, QWK penalizes errors by their ordinal distance from the correct label. As Table 6 shows, even modest QWK improvements indicate meaningful advancements in differentiating CEFR levels, a distinction especially relevant given the skewed ZAEBUC dataset. The significant gains observed in the multi-reference setting with generated GPT-4o essays stem from its flexibility. This evaluation approach treats all three human-assigned CEFR labels as valid references, accommodating natural scoring variations and increasing the chance that model predictions align with at least one reference label. Our analysis revealed that while GPT-4o is powerful for generating diverse content, it struggles to precisely follow the nuanced distribution and specific linguistic features, including error patterns, observed in the manually annotated ZAEBUC dataset. In the GPT-based error injection approach, error type selection is guided by average error counts from the ZAEBUC corpus, but error realization depends on GPT-4o’s interpretation of the prompt, making it less predictable. This inherent challenge in mimicking human-like linguistic and error distributions through zero-shot generation directly contributed to the observed lower agreement rate. In contrast, the controlled method employs a BERT-based classifier for error prediction and applies transformations using bigram-MLE. This systematic approach resulted in a more robust replication of empirically observed error patterns, leading to its superior performance in the average-reference setting. This is expected, as BERT-injected errors more closely resemble learner writing and align more closely with average human ratings. Overall, our findings highlight a trade-off between error alignment and fluency in data augmentation. Controlled error injection excels in the average-reference setting due to its closer alignment with learner errors, while GPT-based augmentation benefits from multi-reference flexibility but less reliably replicates authentic errors. The controlled BERT-based method thus serves as a key component of our pipeline, effectively addressing the limitations of direct GPT error injection.
Qualitative Analysis
The qualitative analysis of the generation process revealed various biases in the GPT-4o outputs, including cultural, gender, and ideological biases. For instance, the essays frequently referenced traditional Arabic themes, reinforced stereotypical gender roles, and reflected culturally narrow assumptions. A clear example of religious bias is that \<الجمعة> ‘Friday’ was selected as the favorite day in all 20 generated essays. Additionally, there was a noticeable tendency to use masculine forms throughout the texts. Such biases may unintentionally disadvantage students whose writing reflects different experiences, perspectives, or identities. Examples of these biases, along with their frequencies, are provided in Appendix A.2. We also observed a lack of diversity among the ten essays generated per prompt, with GPT-4o often repeating similar lexical and structural patterns.
@@ 8 Conclusions and Future Work
This paper presents a hybrid framework for Arabic AES, using LLMs and transformers to tackle data scarcity by generating synthetic essays that partly replicate Arabic learner writing. Building on the ZAEBUC corpus, we developed CEFR-aligned linguistic and error profiles and used GPT-4o to produce 3,040 essays across 152 prompts. However, GPT-4o’s performance relies heavily on prompt engineering, achieving only 27.5% alignment with our reference profiles. To introduce errors, we compare our two methods: (1) GPT-4o prompted multi-step error injection, and (2) our controlled method fine-tuning the CAMeLBERT MSA model to inject errors proportionally to their profiled occurrence. Evaluated with a fine-tuned BERT classifier, our hybrid framework, combining GPT-generated data with controlled error injection, outperformed the baseline (QWK: 27.87%, F1: 38.02% ), offering more reliable and interpretable results. These findings demonstrate the effectiveness of controlled error injection in capturing learner error distributions across CEFR levels. For future work, we will prioritize integrating a human evaluation into our framework. Human annotators will assess the fluency and naturalness of synthetic essays, as well as the realism of injected errors, ensuring that they reflect typical learner patterns at specific CEFR levels. To improve generalizability, we also plan to expand the diversity of prompts beyond predefined topics and incorporate a wider set of writing traits, including coherence, logical flow, and topic relevance, beyond syntactic and lexical features. We also intend to enhance CEFR-level modelling by incorporating more manually annotated essays. This will help capture nuanced linguistic variations across levels and increase the robustness of our dataset. Lastly, we aim to deploy the AES system as an interactive tool to provide users with instant feedback on errors and proficiency levels.
Limitations
@@
Despite the effectiveness of our hybrid Arabic AES framework, we note several limitations related to the quality of generated Arabic essays, error injection accuracy, and the generalization of the AES model. The lack of A1 and C2 essays in ZAEBUC means that there is no gold reference data for these levels, which may impact both linguistic and error profiles, affecting the accuracy of GPT-generated essays. Furthermore, different biases are present in both the ZAEBUC dataset and GPT-4o outputs as discussed in (§7)In addition, due to the lack of comprehensive gold data, GPT struggles to fully replicate real learner writing styles, achieving only 27.5% agreement with linguistic feature profiles. Another limitation is the model’s ability to generalize across various domains and question types. The AES system may struggle with broader writing tasks and alternative prompts since the dataset and augmentation methods focus on predefined prompts. Relying solely on CEFR as a holistic scoring method limits interpretability. Enhancing the dataset with multi-trait annotations, such as coherence, argumentation, and organization, could improve scoring accuracy and feedback quality. Moreover, better-controlled GPT prompting could refine the quality and diversity of generated essays, reducing biases and improving alignment with real learner writing patterns. Due to resource constraints, human evaluation was not feasible in this study; however, we plan to engage CEFR-trained annotators in the future.
Ethical Considerations
@@
Appendix A GPT-4o Generated Essays
A.1 Statistics
CEFR #Essays #Words #Sentences #Tokens Avg_W_L Avg_S_L Unique_Tokens Unique_Words
A1 470 39367 6721 47612 4.66 6.08 4518 4512
A2 470 54980 5565 63769 4.82 10.46 7147 7143
B1 530 100185 7276 113672 4.97 14.62 10742 10734
B2 530 127029 7804 142995 5.03 17.32 12522 12509
C1 520 136849 7630 152962 5.16 19.05 12832 12823
C2 520 146253 7946 163461 5.16 19.57 13686 13676
Total 3040 604663 42942 684471 5.04 14.94 27286 27272
Table 7: Summary statistics of the generated Arabic synthetic essay corpus across CEFR levels. #Essays denotes the number of essays; #Words refers to the total word count; #Sentences indicates the total number of sentences; #Tokens represents the total number of tokens (vocabulary items); Avg_W_L corresponds to the average word length in characters; Avg_S_L refers to the average sentence length in words; lexical diversity is captured through the counts of unique tokens and unique words.
A.2 Bias in the Generated Essays
Arabic Prompt English Prompt Arabic Response English Response Occurrences /20 Bias
\< ما هو يومك المفضل؟ > What is your favorite day? \<الجمعة > Friday 20 Cultural/Religious Bias
\< ماهو طعامك المفضل؟ > What is your favorite food? \< البيتزا > Pizza 16 Globalization Bias
\< ما هي هوايتك المفضلة؟ > What is your favorite hobby? \< القراءة > Reading 10 Socioeconomic/Class Bias
\< رحلة ذهبت إليها > A trip you went on \< ذهبت إلى البحر > I went to the beach 18 Geographical/Cultural Bias
\< ماذا تفعل في عطلة نهاية الأسبوع؟ > What do you do on the weekend? \< نذهب إلى الحديقة > We go to the park 20 Geographical/Cultural Bias
\< ماهي المادة الدراسية المفضلة؟ > What is your favorite school subject? \< الرياضيات > Mathematics 17 Educational System Bias
\< ما هي رياضتك المفضلة؟ > What is your favorite sport? \< كرة القدم > Football 20 Cultural Bias
\< شخص تعتبره مثلك الأعلى > A person you consider your role model \< والدي > My father 17 Gender Bias
\< من هو أفضل صديق لك؟ > Who is your best friend? \< أحمد > Ahmed 20 Gender/Name Bias
\< المهنة المستقبلية > Your future profession \< طبيب > Doctor 13 Stereotype Bias
\< معلم شهير في العالم الأدبي > A famous teacher in the world of literature \< الأهرامات > The Pyramids 17 Cultural Bias
\< أفضل فصول السنة > Your favorite season of the year \< الصيف > Summer 20 Climate Bias
\< لغة تود تعلمها > A language you would like to learn \<الأسبانية > Spanish 16 Language Bias
\< بلد ترغب في السفر إليه > A country you would like to visit \< مصر > Egypt 10 National Identity Bias
\< قوة خارقة تتمناها > A superpower you wish to have \< الطيران > Flying 19 Media Bias
Table 8: Examples of response biases in GPT-4o generated essays.
Appendix B GPT-4o Error Injection Algorithm
Appendix C Error Types Taxonomy
11-Classes 42-Classes Error Description Correct Word Erroneous Word
Morphology (M) M MI Inflection \< عارف> \< معروف>
MT Tense \< ذهب> \< يذهب>
Orthography (O) O OA Alef-Maqsura \< القاضي> \< القاضى >
OA+OH Alef-Maqsura + Hamza \< أضحى > \< اضحا >
OA+OR Alef-Maqsura + Wrong Character \< كشيء> \< كشىء>
OC Chatacter Order \< المدرسة > \< المردسة>
OD Extra Character \< هذا> \< هاذا>
OD+OG Extra Character + Lengthening Short Vowels \< تتطورو > \< تتطور >
OD+OH Extra Character + Hamza \< لأنهم > \< ألانهم >
OD+OM Extra Character + Missing Character \< الاجتماعي > \< الاجتاعيي >
OD+OR Extra Character + Wrong Character \< الصور > \< السوور>
OH Hamza \< العب > \< إلعب >
OH+OM Hamza + Missing Character \< الأشياء > \< الاشيا >
OH+OT Hamza + Ta-Marbuta \< إمارة > \< اماره >
OM Missing Character \< المدرسة > \< المدسة >
OM+OR Missing Character + Wrong Character \< المجتمع > \< الجطمع >
OR Wrong Character \< المدرسة> \< المدرصة >
OR+OT Wrong Character + Ta-Marbuta \<مكتظة> \< مكتضه>
OT Ta-Marbuta \< غرفة> \< غرفه >
OW Alef-Fariqa \< كتبوا > \< كتبو >
Semantics (S) S SF Conjunction \< فسبحان > \< سبحان >
SW Word Selection \< على > \< من >
Punctuation
P P Punctuation \<السوق،> \<السوق.>
Syntax (X) X XC Case \<رائعا> \<رائع>
XC+XG Case + Gender \<مجتهدا> \<مجتهدة>
XC+XN Case + Number \<نواح> \<نواحي>
XF Definiteness \<المفيد> \<مفيد>
XG Gender \<كان> \<كانت>
XM Missing Word \<على> NULL
XN Number \<كتابين> \<كتب>
XT Unnecessary Word NULL \<على>
Combination (Comb.) M+O MI+OH Inflection + Hamza \< أشخاص > \< اشخاصك >
O+X OH+XC Hamza + Case \< أضرارا> \< اضرار >
SPLIT SPLIT SPLIT Split \<دولة الإمارات> \<دولةالإمارات>
MERGE MERGE MERGE Merge \< بالعلم > \< ب العلم >
DELETE DELETE DELETE Delete NULL \<داخل>
Table 9: Illustrative examples of error types categorized according to the ARETA error taxonomy Belkebir and Habash (2021). The table presents hierarchical mappings from coarse-grained (11-Class) to fine-grained (42-Class) error categories, alongside representative corrections.